# Towards Diverse and Efficient Audio Captioning via Diffusion Models

### v1

@@ IV-B Diversity and Efficiency
We present the diversity and efficiency evaluation results in table II, where we also report the model parameter size. We demonstrate that DAC models achieve significant diversity at both the lexical and sentence levels, while maintaining a relatively small model size and fast generation speed. In terms of diversity, only the Qwen-audio series surpass DAC, but with a much larger parameter size and a much slower inference speed. Audio-Flamingo outperforms ACT and HTSAT-BART with its pre-trained flamingo backbone, but lags behind the DAC models, particularly in terms of sentence diversity. We note that a stronger LLM backbone can indeed help increase generation diversity and quality, but this comes at the cost of greater computational resources and slower generation speed. Regarding generation efficiency, the fastest model, ACT, is almost eight times faster than the slowest Audio Flamingo with the same computational resources. Our DAC models are also significantly faster than AR models while maintaining competitive performance.
@@
Beyond the quantitative results, we also observe two advantages that may not be evident from the table. Firstly, DAC is not sensitive to caption length due to their holistic denoising and decoding process on embeddings, as long as the length does not exceed the model’s token limit. This benefits both the generation speed and overall generation quality. For DAC, the generation token length limit is set to 40, while the average caption length for Audiocaps is around 10. In contrast, AR models experience a linear increase in generation time with sequence length. Longer sequences may also lead to distraction challenges in LLMs. Secondly, a smaller parameter size facilitates easier deployment on other devices or allows for larger batch sizes. Models with LLMs as backbones are challenging to deploy on most devices with billion-level parameter size.
@@
Figure 2: Change of the metric scores on the Audiocaps dataset during fine-funing following a pre-training on the Audioset dataset. We demonstrate that achieving good convergence during Audioset pre-training does not necessarily translate to a strong initial performance on the AudioCaps test set, especially for metrics like BLEU and SPIDEr.
IV-C Further Discussion about the Metrics
@@
An intriguing observation during the DAC fine-tuning process on Audiocaps is that achieving good convergence during the pre-training phase, e.g., pre-training on Audioset, does not guarantee a strong initial performance on the AudioCaps test set for certain metrics, as illustrated in fig. 2. Although pre-training on a significantly larger dataset yields much better results from the loss perspective (around 5 for Audiosat vs around 7 for AudioCaps), metrics such as Bleu, CIDEr, and SPIDEr are considerably lower than the baseline level if we directly evaluate the checkpoint on the Audiocaps dataset. In contrast, metrics like CLAP and BERT maintain consistently.
@@
We thus consider metrics such as CLAP and BERT to be more equitable and convincing scores for evaluating captioning ability in our experiment. Traditional metrics like Bleu and CIDEr are based on token-level matching, implying that a better imitation of the ground truth strings results in a higher score. However, our findings suggest that a good caption does not necessarily have to replicate the ground truth verbatim. As demonstrated in fig. 2, there are instances where the generated captions are semantically similar to both the audio and the ground truth text, yet they receive low scores due to differences in caption style or grammar. Fine-tuning within the target distribution also does not necessarily enhance overall semantic similarity. On the other hand, metrics like CLAP and BERT measure the overall semantic similarity of the generated captions and the ground truth. On these two metrics, our model outperforms the other baselines.
V Conclusion
@@
In this paper, we present Diffusion-based Audio Captioning (DAC), a state-of-the-art diffusion-based model for generating diverse and high-quality audio captioning. Leveraging the NAR architecture, DAC excels in providing a superior mix of diversity, efficiency, and quality in generating textual descriptions of audio pieces, outperforming existing AR models. Our model surpasses current benchmarks, especially in terms of generation diversity and processing speed, while maintaining a lightweight design. We hope that our findings will inspire further exploration of a diffusion-based comprehensive framework for multimodal content generation.

### v2

@@ 4.2 Diversity and Efficiency
We present the diversity and efficiency evaluation results in table 2, where we also report the model parameter size. We demonstrate that DAC models achieve significant diversity at both the lexical and sentence levels, while maintaining a relatively small model size and fast generation speed. In terms of diversity, only the Qwen-audio series surpass DAC, but with a much larger parameter size and a much slower inference speed. A stronger LLM backbone can indeed help increase generation diversity and quality, but this comes at the cost of greater computational resources and slower generation speed. Regarding generation efficiency, the fastest model, ACT, is almost eight times faster than the slowest Audio Flamingo with the same computational resources. Our DAC models are also significantly faster than AR models while maintaining competitive performance. Compared to DAC-RLD, our model is competitive in generation speed but outperforms it in diversity. We attribute this to the DAC-RLD design, where the subsequent AR decoder from the BART module may mask the potential of diffusion in enhancing diversity.
@@
Beyond the quantitative results, we observe two advantages that may not be evident from the table. Firstly, DAC is not sensitive to caption length due to their holistic denoising and decoding process on embeddings, as long as the length does not exceed the model’s token limit. This benefits both the generation speed and overall generation quality. For DAC, the generation token length is limited to 40, while the average caption length for AudioCaps is around 10. In contrast, AR models experience a linear increase in generation time with sequence length. Longer sequences may also lead to distraction challenges in LLMs. Secondly, a smaller parameter size facilitates easier deployment on devices or allows for larger batch sizes.
@@
Figure 2: Change of the metric scores on the AudioCaps dataset during fine-funing following pre-training. We demonstrate that achieving good convergence during Audioset pre-training does not necessarily translate to a strong initial performance on the AudioCaps test set, especially for metrics like BLEU and SPIDEr.
4.3 Further Discussion about the Metrics
@@
An intriguing observation during the DAC fine-tuning process on AudioCaps is that achieving good convergence during the pre-training phase, e.g., pre-training on Audioset, does not guarantee a strong initial performance on the AudioCaps test set for certain metrics, as illustrated in fig. 2. Although pre-training on a significantly larger dataset yields much better results from the loss perspective (around 5 for Audioset vs around 7 for AudioCaps), metrics such as BLEU, CIDEr, and SPIDEr are considerably lower than the baseline level if we directly evaluate the checkpoint on the AudioCaps dataset. In contrast, metrics like CLAP and BERT are maintained consistently.
@@
We thus consider metrics such as CLAP and BERT to be more equitable and convincing scores for evaluating captioning ability in our experiment. Traditional metrics like BLEU and CIDEr are based on token-level matching, implying that a better imitation of the ground truth strings results in a higher score. However, our findings suggest that a good caption does not necessarily have to replicate the ground truth verbatim. As demonstrated in fig. 2, there are instances where the generated captions are semantically similar to both the audio and the ground truth text, yet they receive low scores due to differences in caption style or grammar. Fine-tuning within the target distribution also does not necessarily enhance overall semantic similarity. On the other hand, metrics like CLAP and BERT measure the overall semantic similarity. On these two metrics, our model outperforms the other baselines.
5 Conclusion
@@
We present DAC, a diffusion-based NAR model for generating diverse and high-quality audio captioning. Leveraging the NAR architecture, DAC excels in providing a superior mix of diversity, efficiency, and quality in generating textual descriptions of audio pieces, outperforming several existing AR models. Our model surpasses current benchmarks, especially in terms of generation diversity and processing speed, while maintaining a lightweight design. We hope that our findings will inspire further exploration of a diffusion-based comprehensive framework for multimodal content generation.